OpenStudy (rock_mit182):

Simple Equation Involving x and y : \[\frac{y-1}{ x+3 } =-\frac{ 3 }{ 4 }\]

OpenStudy (rock_mit182):

@wolf1728

OpenStudy (rock_mit182):

someone please explain step by step

OpenStudy (nincompoop):

well now.. solve them independently

OpenStudy (nincompoop):

what is y - 1 = 3

OpenStudy (nincompoop):

and what is x+1 = 4

OpenStudy (rock_mit182):

I mean, can i just do the cross muliplying ?

OpenStudy (nincompoop):

why would you cross multiply? you're not doing a proportion you're solving for possible values of x and y

OpenStudy (rock_mit182):

\[(y-1)*4=-3*(x+3)\]

OpenStudy (nincompoop):

go ahead

OpenStudy (rock_mit182):

:/ mmm so...

OpenStudy (nincompoop):

that should give you a distributive property on the left and on the right side of the equation try solving it with your method

OpenStudy (rock_mit182):

i guess is wrong already, what do i have to do, please i can't fail on this one

OpenStudy (nincompoop):

do you have ten fingers?

OpenStudy (rock_mit182):

ahh ? i guess is a JK thing of yours

OpenStudy (nincompoop):

I want to know if you have ten fingers, because I don't want to assume that you do

OpenStudy (rock_mit182):

hey dude, i doin on't have a lot of time like you think so, so I'm going to solve in the way i told you

OpenStudy (nincompoop):

okay then do it your way I was going to show you a trick that you can base your future algebra problems with

OpenStudy (nincompoop):

if you don't have ten fingers just tell me, I won't judge you

OpenStudy (rock_mit182):

\[4y-4 = -3x-9\]

OpenStudy (nincompoop):

did it help you solve for x and y?

OpenStudy (nincompoop):

you're doing an equation, where is that solution going to lead you?

OpenStudy (rock_mit182):

to nowhere

OpenStudy (nincompoop):

exactly

OpenStudy (nincompoop):

we're going to try to solve for y put three fingers up

OpenStudy (rock_mit182):

maybe a function thing..

OpenStudy (nincompoop):

just put three fingers up

OpenStudy (rock_mit182):

lol go ahead

OpenStudy (nincompoop):

okay it says y - 1 = 3 with your three fingers up, that is your right hand side of the equation now on the left hand side it says y - 1 how about you add 1 to find out how many you had before one was deducted

OpenStudy (nincompoop):

done?

OpenStudy (nincompoop):

now you know how the mechanism is done, let me lay it out for you y - 1 = 3 //add 1 to both sides of the equation y -1 + 1 = 3 + 1 //the left hand side becomes y because -1 + 1 becomes zero y + 0 = 4 // finalize y = 4 do the same thing with the x in the denominator and tell me what you got

OpenStudy (rock_mit182):

i have to describe every step

OpenStudy (rock_mit182):

what kind of property is that ?

OpenStudy (rock_mit182):

i need t

OpenStudy (rock_mit182):

to inform! please come back dude. please help me out

OpenStudy (nincompoop):

it applies to variables too

OpenStudy (rock_mit182):

I know that ! what i was saying is related with the whole approach in the equation

OpenStudy (nincompoop):

it has no specific name like cross multiplication

OpenStudy (rock_mit182):

I'm going to trust in you,but if i fail, iLL be screwed..

OpenStudy (nincompoop):

you won't how about you tell me what you got for x?

OpenStudy (nincompoop):

because we are not done yet

OpenStudy (rock_mit182):

x=1

OpenStudy (nincompoop):

okay plug those into the original equation but you will notice that the right hand side is a negative

OpenStudy (rock_mit182):

yes...

OpenStudy (nincompoop):

the rule is that for a fraction to be negative, either the numerator or denominator has to be negative

OpenStudy (nincompoop):

so let us tweak either the numerator or denominator. you pick which would you like to be the negative on the right hand side

OpenStudy (rock_mit182):

I didn't know that i can solve the equation in this way, hope is right

OpenStudy (rock_mit182):

y/x = -4 ?

OpenStudy (rock_mit182):

doesn't make sense for me

OpenStudy (rock_mit182):

\[\frac{y-1}{ x+3 } =-\frac{ 3 }{ 4 }\]

OpenStudy (rock_mit182):

4y=−3x−5, \[y =-\frac{ 3x }{ 4 }-\frac{ 5 }{ 4 }\]

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

just pick if you want the numerator or the denominator on the right side to be negative

OpenStudy (rock_mit182):

dude thanks anyway...

OpenStudy (nincompoop):

okay